# Cosmic (super)strings with a time-varying tension

### v1

@@ 1 Introduction
Cosmic strings Kibble:1976sj ; Kibble:1984hp ; Hindmarsh:1994re ; Vilenkin:2000jqa ; Copeland:2009ga are among the most promising signatures of high-energy physics which could still be present in the late universe. In field theory, these strings are solitonic remnants of a broken global or gauge symmetry, with a tension $\mu$ given by the energy scale of symmetry breaking. In String Theory, they arise as fundamental objects, and their tension is related to the new fundamental scale of the theory. While initially considered as promising candidates for the seed of structure formation, the dimensionless combination $G\mu$ (where $G$ is Newton’s constant) is now too severely constrained by Cosmic Microwave Background (CMB) Charnock:2016nzm ($G\mu\lesssim 10^{-7}$) and Pulsar Timing Arrays  EPTA:2023xxk ; NANOGrav:2020bcs ; NANOGrav:2023gor ($G\mu\lesssim 10^{-10}$) measurements. However, it is important to emphasize that the study of cosmic strings remains a very active topic because of their potential to seed dark matter as the QCD axion (see Gorghetto:2018myk ; Gorghetto:2020qws and references therein), gravitational radiation (see the recent reviews Caprini:2018mtu ; Auclair:2019wcv ; Sousa:2024ytl ) and other phenomenological features for which we refer to Vilenkin:2000jqa .
@@
Regarding gravitational waves (GWs), it is particularly interesting that the strings are known to enter a scaling regime [^1] Strictly speaking, a departure from the scaling regime occurs when the equation of state of the background changes Sousa:2013aaa ; Gouttenoire:2019kij , but we will restrict ourselves to periods with constant equation of state. In addition, we will consider scaling loosely and not take into account possible logarithmic violations due to the IR divergence of the tension of global $U(1)$ strings which is a source of disagreement in the literature, see Correia:2024cpk ; Saikawa:2024bta . We thank Amelia Drew for discussions on this point., for which a network of $\mathcal{O}(1)-\mathcal{O}(10)$ super-Hubble cosmic strings contribute a constant fraction to the energy density of the background. The scaling regime is achieved due to a balance of the energy density that enters the horizon as the Universe expands and the energy loss of the strings into small loops, which then decay predominantly [^2] How much energy is deposited in massless radiation is another source of disagreement in the literature, see e.g. Hindmarsh:2021mnl . Our main conclusions regarding scaling should not be altered by the decay channel of the loops, but the discussion on loop dynamics of section 4.3 might have to be reconsidered in the future. to axions or gravitational waves. Our interest in this article is the latter case, which is expected for local strings and superstrings, further described below. Due to the scaling properties of the string configuration, the GWs sourced at every e-fold contribute a meaningful fraction to the amplitude of the GW background today, with a spectral index that depends on the equation of state of the Universe at the time the GWs are sourced Cui:2017ufi ; Cui:2018rwi (the emission time can be inferred from the frequency of the GWs today). The GW spectrum from a network of cosmic strings therefore provides a map of the cosmological history since their formation: a golden opportunity to test the cosmology of the very early Universe.
@@
On the other hand, one of the most generic consequences of String Theory, existence of fundamental strings aside, is the fact that there exist no free parameters, with all couplings and dimensionful quantities arising from expectation values of scalar fields - moduli - and the fundamental string scale $\sqrt{\alpha^{\prime}}$. It follows that the properties of cosmic strings, whether field theoretic or fundamental, generically depend on the moduli in a stringy setting. These scalar fields have flat potentials at tree level and their vacuum expectation values (VEVs) parametrise degeneracies in the space of solutions to the tree level equations describing the low-energy vacua of string theory. Higher order corrections generically lift these degeneracies and result in a potential and a mass scale for the moduli, typically smaller than the string scale (see McAllister:2023vgy for a recent review of the problem of moduli stabilization). It is therefore sensible to consider scenarios where the moduli are evolving on cosmological timescales, and this will invariably lead to variation of fundamental constants in the early universe, including the tension of cosmic string networks. The GW spectrum arising from these strings will therefore be modified, providing a potential testing ground for cosmologies featuring time-varying moduli in the early Universe (see Cicoli:2023opf for a review). In order to make concrete statements about the GW spectrum from these strings, it is important to argue that a scaling regime is reached also when the tension varies.
@@
In this paper we explore the phenomenology of cosmic strings with varying tension, which has already received significant attention in Yamaguchi:2005gp ; Ichikawa:2006rw ; Cheng:2008ma ; Wang:2012naa ; Martins:2018dqg ; Emond:2021vts ; Conlon:2024uob . Focusing on power-law cosmologies (with scale factor $a\sim t^{2/n}$) and a polynomially decreasing tension $\mu\sim t^{-m}$, [^3] The $t$ denotes coordinate time in an FLRW cosmology. we provide a comprehensive study of both the long string network and the evolution of loops, correcting some earlier claims in the literature on the former. In particular, we incorporate a time-dependent tension $\mu(t)$ in the VOS model Martins:1996jp ; Martins:2000cs , and show how new attractor solutions can in principle arise alongside the usual scaling regime. However, only the scaling regime can be reached for realistic values of the parameters and/or the initial conditions. To further verify our conclusions, we analytically test some assumptions on the small-scale structure of the strings by solving the equations of motion for a circular loop and for the perturbations of an infinite, static string. We plan on exploiting these results to compute the GW spectrum arising from a network of cosmic superstrings with varying tension in future work us:2024? .
@@
The paper is structured as follows. In Section 2, we will briefly illustrate cosmic strings from the perspective of String Theory. We will provide simple examples of how the tension of the strings generically depends on the moduli of the compactification, both for superstrings and ordinary, field theory strings. In Section 3 we provide a brief summary of the dynamics of Nambu-Goto strings with constant tension, reviewing the Velocity One Scale model (VOS) Martins:1996jp ; Martins:2000cs and the dynamics of averaged loops. Section 4 contains the main results of the paper. After writing down the microscopic equations of motion with varying tension Emond:2021vts ; Conlon:2024uob , we provide an approximate solution for the dynamics of the circular loop, extending on Conlon:2024uob . We appropriately generalize the VOS model and use a dynamical system analysis to argue that scaling attractor points generically exist, provided there is a condition satisfied for the so-called momentum parameter. We find suggestive evidence that this condition is satisfied by exploring the dynamics of small perturbations to the straight, static string. Lastly, we solve the averaged loop equations and comment on the stability of the network under percolation Conlon:2024uob . Section 5 contains our main conclusions and we provide additional details on the physics of dimensional reduction (key to understand moduli dependence in the examples we provide) in Appendix A, the calculation of the solution for the circular loop in Appendix B and further details of the dynamical system analysis in Appendices C and D.
@@ 2 Cosmic strings in String Theory
$m_{s}=\frac{g_{s}^{1/4}M_{P}}{\sqrt{4\pi\mathcal{V}}},$ (1)
where $\mathcal{V}$ is the (Einstein-frame) compactification volume in string units of $\left(2\pi\sqrt{\alpha^{\prime}}\right)^{6}$. One of the main points of this work is that in String Theory all couplings are dynamical, as they depend on the expectation values of scalar fields (the moduli), which can themselves vary with time. In the above example, the latter can be identified with the volume modulus and the dilaton, which sets the string coupling $g_{s}$, but more exotic possibilites can also occur. This motivates us to study the dynamics of cosmic strings with a time-dependent tension, whose rich phenomenology will be analysed in the following chapters. However, we also emphasize how the results that we obtain in the rest of the paper may be used to study field theory strings with a time dependent tension, which can equally arise within string theory compactifications. In the following, we discuss both cases in detail, as well as providing some explicit examples. This section is only meant as a general motivation towards the rest of the analysis, and the details are not important to follow the rest of the paper. The uninterested reader may skip directly to section 3.
@@ 2.1 Cosmic superstrings
The study of cosmic superstrings has a rich and varied history, beginning with Witten:1985fp ; Polchinski:1988cn ; Copeland:2003bj ; Damour:2004kw ; Jackson:2004zg ; Copeland:2009ga . At the microscopic level, they can be realised as either fundamental (F-)strings, or solitonic $D1$-branes (D-strings), which are dual to F-strings under S-duality. [^5] The combination of $p$ F- and $q$ D-strings can also form stable bound states known as $(p,q)$-strings, see Witten:1985fp ; Polchinski:1988cn ; Copeland:2003bj ; Damour:2004kw ; Jackson:2004zg ; Copeland:2009ga . Both types can be easily produced by one of the most popular inflationary mechanisms in String Theory, $D3-\overline{D3}$ brane inflation [^6] One can also obtain cosmic strings in 4 dimensions from a $p$-brane wrapped around a $p-1$ compact cycle. While such defects can be produced in generic brane-antibrane inflation, it is not the case for the most developed scenario of $D3-\overline{D3}$ annihilation. Dvali:1998pa ; Burgess:2001fx ; Jones:2002cv ; Sarangi:2002yt ; Jones:2003da ; Pogosian:2003mz ; Kachru:2003sx ; Burgess:2022nbx ; Cicoli:2024bwq , or more generally when the energy density of the Universe becomes order one in string units Aharonov:1987ah ; Polchinski:2004ia   [^7] Whether the strings percolate and form a network of long, cosmic superstrings or reach thermal equilibrium instead depends on the details of the system, as explained in Abel:1999rq ; Frey:2023khe .. In this section, we will briefly summarize the main properties of F- and D-strings (tension, re-connection probability, etc) relevant to their cosmological evolution.
To begin with, a crucial requirement for super-strings to be cosmologically viable is a tension (string scale) that lies many orders of magnitude below the Planck scale. Indeed, even the most conservative bounds arising from CMB observations give rise to the hard limit $G\mu<10^{-7}$ Charnock:2016nzm , while NANOGrav data NANOGrav:2020bcs ; NANOGrav:2023gor points to even lower values of $G\mu\lesssim 10^{-11}-10^{-12}$ Blanco-Pillado:2021ygr ; Ellis:2023tsl . In heterotic string compactifications, for example, the string scale is generically of order $G\mu\sim\frac{\alpha_{\rm{GUT}}}{16\pi}\sim 10^{-3}$, and cosmic strings are immediately ruled out Witten:1985fp . [^8] Moreover, various instabilities are generically present, so that a stable network cannot be formed. To lower the string tension down to acceptable values, two main classes of mechanisms have been proposed, mainly in the context of type IIB compactifications: large extra dimensions and warping (or a combination thereof). In the following, we will review some their key properties, as well as those of the resulting cosmic superstrings.
@@ 2.1.1 Large extra dimensions
The first possibility is to consider a scenario where the extra dimensions are large in string units, and the string scale is diluted by the volume of the compactification manifold. Scenarios with large extra dimensions first appeared as a phenomenological solution to the hierarchy problem Arkani-Hamed:1998jmv , and there is a natural string theory incarnation in terms of the Large Volume Scenario (LVS) Balasubramanian:2005zx ; Conlon:2005ki . The LVS model is a scenario where all moduli can be stabilised at exponentially large values of the volume $\mathcal{V}$, allowing a very natural realisation of hierarchies between different scales. As an example, one can easily see from (1) that exponentially large volumes can introduce a significant separation between the the 4d Planck scale and the string scale, allowing one to bypass the bounds discussed in the previous section on the tension of cosmic strings. Realistic values of the stabilised volume in LVS are of order $\mathcal{V}\sim 10^{6}-10^{13}$ (and $g_{s}\sim 0.1)$, corresponding (from (1)) to a string tension in the range $G\mu\sim 10^{-9}-10^{-16}$, which can easily evade observational bounds.
@@
However, as emphasized in Conlon:2022pnx , this need not be the case during the entire history of the universe, as such constraints only apply from the CMB emission time onwards. It is indeed possible to have a string scale that is marginally lower then the 4d Planck scale in the very early universe (say immediately after inflation), which is subsequently lowered by the rolling of the volume modulus along its potential. Given the observed hierarchy between the electroweak and inflationary scales, $\Lambda_{{\rm EW}}\ll\Lambda_{{\rm inf}}$, this is not only possible but actually a very desirable feature of a realistic model. In the context of the LVS scenario such a cosmology has been studied in detail Conlon:2022pnx ; Apers:2022cyl ; Revello:2023hro ; Apers:2024ffe ; Conlon:2024hgw ; Conlon:2024ene ; Apers:2024dtn , consisting in a succession of different kination and tracker epochs mediated by the volume (see also the earlier work Conlon:2008cj ).
@@
The presence of large extra dimensions also reduces the probability for strings to scatter off each other, as they can now miss each other in the extra dimensions. If the strings were truly able to propagate within the extra-dimensional bulk, this would lower the reconnection probability by a factor of the volume, and quickly reach very small values. However, the position of the strings along the extra dimensions is a worldsheet scalar and corrections to e.g. the flat space metric will stabilize it. Using this logic, it has been argued Jackson:2004zg that warping effects (generically arising due to the presence of sources of energy along the extra dimensions) localize the strings in a small extra-dimensional region, yielding a reconnection probability of
$P\simeq 0.5g_{s}^{2}\left[\frac{6\pi}{\log\left(\frac{(2\pi)^{7}\mathcal{V}}{\alpha^{\prime 3}}\right)}\right],$ (2)
@@
which depends very weakly on the volume. It is worth noting that brane-antibrane inflation, which naturally produces cosmic superstrings at its endpoint, has recently been reasessed Burgess:2022nbx ; Cicoli:2024bwq and a potential with a hierarchical separation between the inflationary and late-time scales has been obtained. The volume modulus would evolve from its value during inflation to its exponentially large value at the true, late-time minimum, providing a concrete example of the ideas motivating this article.

### v2

@@ 1 Introduction
Cosmic strings [1, 2, 3, 4, 5] are among the most promising signatures of high-energy physics which could still be present in the late universe. In field theory, these strings are solitonic remnants of a broken global or gauge symmetry, with a tension $\mu$ given by the energy scale of symmetry breaking. In String Theory, they arise as fundamental objects, and their tension is related to the new fundamental scale of the theory. While initially considered as promising candidates for the seed of structure formation, the dimensionless combination $G\mu$ (where $G$ is Newton’s constant) is now too severely constrained by Cosmic Microwave Background (CMB) [6] ($G\mu\lesssim 10^{-7}$) and Pulsar Timing Arrays  [7, 8, 9, 10] ($G\mu\lesssim 10^{-10}$) measurements. However, it is important to emphasize that the study of cosmic strings remains a very active topic because of their potential to seed dark matter as the QCD axion (see [11, 12] and references therein), gravitational radiation (see the recent reviews [13, 14, 15]) and other phenomenological features for which we refer to [4].
@@
Regarding gravitational waves (GWs), it is particularly interesting that the strings are known to enter a scaling regime [^1] Strictly speaking, a departure from the scaling regime occurs when the equation of state of the background changes [16, 17], but we will restrict ourselves to periods with constant equation of state. In addition, we will consider scaling loosely and not take into account possible logarithmic violations due to the IR divergence of the tension of global $U(1)$ strings which is a source of disagreement in the literature, see [18, 19]. We thank Amelia Drew for discussions on this point., for which a network of $\mathcal{O}(1)-\mathcal{O}(10)$ super-Hubble cosmic strings contribute a constant fraction to the energy density of the background. The scaling regime is achieved due to a balance of the energy density that enters the horizon as the Universe expands and the energy loss of the strings into small loops, which then decay predominantly [^2] How much energy is deposited in massless radiation is another source of disagreement in the literature, see e.g. [20, 21, 22]. Our main conclusions regarding scaling should not be altered by the decay channel of the loops, but the discussion on loop dynamics of section 4.3 might have to be reconsidered in the future, and so would the PTA constraints (see [23] for discussions on this point and [24] for opportunities for multi-messenger opportunities). to axions or gravitational waves. Our interest in this article is the latter case, which is expected for local strings and superstrings, further described below. Due to the scaling properties of the string configuration, the GWs sourced at every e-fold contribute a meaningful fraction to the amplitude of the GW background today, with a spectral index that depends on the equation of state of the Universe at the time the GWs are sourced [25, 26] (the emission time can be inferred from the frequency of the GWs today). The GW spectrum from a network of cosmic strings therefore provides a map of the cosmological history since their formation: a golden opportunity to test the cosmology of the very early Universe.
@@
On the other hand, one of the most generic consequences of String Theory, existence of fundamental strings aside, is the fact that there exist no free parameters, with all couplings and dimensionful quantities arising from expectation values of scalar fields - moduli - and the fundamental string scale $\sqrt{\alpha^{\prime}}$. It follows that the properties of cosmic strings, whether field theoretic or fundamental, generically depend on the moduli in a stringy setting. These scalar fields have flat potentials at tree level and their vacuum expectation values (VEVs) parametrise degeneracies in the space of solutions to the tree level equations describing the low-energy vacua of string theory. Higher order corrections generically lift these degeneracies and result in a potential and a mass scale for the moduli, typically smaller than the string scale (see [27] for a recent review of the problem of moduli stabilization). It is therefore sensible to consider scenarios where the moduli are evolving on cosmological timescales, and this will invariably lead to variation of fundamental constants in the early universe, including the tension of cosmic string networks. The GW spectrum arising from these strings will therefore be modified, providing a potential testing ground for cosmologies featuring time-varying moduli in the early Universe (see [28] for a review). In order to make concrete statements about the GW spectrum from these strings, it is important to argue that a scaling regime is reached also when the tension varies.
@@
In this paper we explore the phenomenology of cosmic strings with varying tension, which has already received significant attention in [29, 30, 31, 32, 33, 34, 35]. Focusing on power-law cosmologies (with scale factor $a\sim t^{2/n}$) and a polynomially decreasing tension $\mu\sim t^{-m}$, [^3] The $t$ denotes coordinate time in an FLRW cosmology. we provide a comprehensive study of both the long string network and the evolution of loops, correcting some earlier claims in the literature on the former. In particular, we incorporate a time-dependent tension $\mu(t)$ in the VOS model [36, 37], and show how new attractor solutions can in principle arise alongside the usual scaling regime. However, only the scaling regime can be reached for realistic values of the parameters and/or the initial conditions. To further verify our conclusions, we analytically test some assumptions on the small-scale structure of the strings by solving the equations of motion for a circular loop and for the perturbations of an infinite, static string. We plan on exploiting these results to compute the GW spectrum arising from a network of cosmic superstrings with varying tension in future work [38].
@@
The paper is structured as follows. In Section 2, we will briefly illustrate cosmic strings from the perspective of String Theory. We will provide simple examples of how the tension of the strings generically depends on the moduli of the compactification, both for superstrings and ordinary, field theory strings. In Section 3 we provide a brief summary of the dynamics of Nambu-Goto strings with constant tension, reviewing the Velocity One Scale model (VOS) [36, 37] and the dynamics of averaged loops. Section 4 contains the main results of the paper. After writing down the microscopic equations of motion with varying tension [34, 35], we provide an approximate solution for the dynamics of the circular loop, extending on [35]. We appropriately generalize the VOS model and use a dynamical system analysis to argue that scaling attractor points generically exist, provided there is a condition satisfied for the so-called momentum parameter. We find suggestive evidence that this condition is satisfied by exploring the dynamics of small perturbations to the straight, static string. Lastly, we solve the averaged loop equations and comment on the stability of the network under percolation [35]. Section 5 contains our main conclusions and we provide additional details on the physics of dimensional reduction (key to understand moduli dependence in the examples we provide) in Appendix A, the calculation of the solution for the circular loop in Appendix B and further details of the dynamical system analysis in Appendices C and D.
@@ 2 Cosmic strings in String Theory
$\mu\equiv m_{s}^{2}=\frac{g_{s}^{1/2}M_{P}^{2}}{4\pi\mathcal{V}},$ (2.1)
where $g_{s}$ is the string coupling, and $\mathcal{V}$ is the (Einstein-frame) compactification volume in string units of $\left(2\pi\sqrt{\alpha^{\prime}}\right)^{6}$. One of the main points of this work is that in String Theory all couplings are dynamical, as they depend on the expectation values of scalar fields (the moduli), which can themselves vary with time. In the above example, the latter can be identified with the volume modulus and the dilaton, which sets the string coupling, but more exotic possibilities can also occur. This motivates us to study the dynamics of cosmic strings with a time-dependent tension, whose rich phenomenology will be analysed in the following chapters. However, we also emphasize how the results that we obtain in the rest of the paper may be used to study field theory strings with a time dependent tension, which can equally arise within string theory compactifications. In the following, we discuss both cases in detail, as well as providing some explicit examples. This section is only meant as a general motivation towards the rest of the analysis, and the details are not important to follow the rest of the paper. The uninterested reader may skip directly to section 3.
@@ 2.1 Cosmic superstrings
The study of cosmic superstrings has a rich and varied history, beginning with [39, 40, 41, 42, 43, 5]. At the microscopic level, they can be realised as either fundamental (F-)strings, or solitonic $D1$-branes (D-strings), which are dual to F-strings under S-duality. [^5] The combination of $p$ F- and $q$ D-strings can also form stable bound states known as $(p,q)$-strings, see [39, 40, 41, 42, 43, 5]. Both types can be easily produced by one of the most popular inflationary mechanisms in String Theory, $D3-\overline{D3}$ brane inflation [^6] One can also obtain cosmic strings in 4 dimensions from a $p$-brane wrapped around a $p-1$ compact cycle. While such defects can be produced in generic brane-antibrane inflation, it is not the case for the most developed scenario of $D3-\overline{D3}$ annihilation. [44, 45, 46, 47, 48, 49, 50, 51, 52], or more generally when the energy density of the Universe becomes order one in string units [53, 54]  [^7] Whether the strings percolate and form a network of long, cosmic superstrings or reach thermal equilibrium instead depends on the details of the system, as explained in [55, 56].. In this section, we will briefly summarize the main properties of F- and D-strings (tension, re-connection probability, etc) relevant to their cosmological evolution.
To begin with, a crucial requirement for super-strings to be cosmologically viable is a tension (string scale) that lies many orders of magnitude below the Planck scale. Indeed, even the most conservative bounds arising from CMB observations give rise to the hard limit $G\mu<10^{-7}$ [6], while Pulsar Timing Array (PTA) data (for example NANOGrav [8, 9, 10]) points to even lower values of $G\mu\lesssim 10^{-11}-10^{-12}$ [57, 58]. In heterotic string compactifications, for example, the string scale is generically of order $G\mu\sim\frac{\alpha_{\rm{GUT}}}{16\pi}\sim 10^{-3}$, and cosmic strings are immediately ruled out [39]. [^8] Moreover, various instabilities are generically present, so that a stable network cannot be formed. To lower the string tension down to acceptable values, two main classes of mechanisms have been proposed, mainly in the context of type IIB compactifications: large extra dimensions and warping (or a combination thereof). In the following, we will review some their key properties, as well as those of the resulting cosmic superstrings.
@@ 2.1.1 Large extra dimensions
The first possibility is to consider a scenario where the extra dimensions are large in string units, and the string scale is diluted by the volume of the compactification manifold. Scenarios with large extra dimensions first appeared as a phenomenological solution to the hierarchy problem [59], and there is a natural string theory incarnation in terms of the Large Volume Scenario (LVS) [60, 61]. The LVS model is a scenario where all moduli can be stabilised at exponentially large values of the volume $\mathcal{V}$, allowing a very natural realisation of hierarchies between different scales. As an example, one can easily see from (2.1) that exponentially large volumes can introduce a significant separation between the 4d Planck scale and the string scale, allowing one to bypass the bounds discussed in the previous section on the tension of cosmic strings. Realistic values of the stabilised volume in LVS are of order $\mathcal{V}\sim 10^{6}-10^{13}$ (and $g_{s}\sim 0.1)$, corresponding (from (2.1)) to a string tension in the range $G\mu\sim 10^{-9}-10^{-16}$, which can easily evade observational bounds.
@@
However, as emphasized in [62], this need not be the case during the entire history of the universe, as such constraints only apply from the CMB emission time onwards. It is indeed possible to have a string scale that is marginally lower then the 4d Planck scale in the very early universe (say immediately after inflation), which is subsequently lowered by the rolling of the volume modulus along its potential. [^9] In principle, the dilaton could also be rolling along its potential and give rise to a decreasing tension. However, the most commonly studied scenarios of moduli stabilization (such as LVS and KKLT) are based on the GKP construction [63], where the axio-dilaton is stabilized at much higher scales together with the complex structure sector. Given the presumed hierarchy between the electroweak and inflationary scales, $\Lambda_{{\rm EW}}\ll\Lambda_{{\rm inf}}$, this is not only possible but actually a very desirable feature of a realistic model. In the context of the LVS scenario such a cosmology has been studied in detail [62, 64, 65, 66, 67, 68, 69], consisting in a succession of different kination and tracker epochs mediated by the volume (see also the earlier work [70]).
@@
The presence of large extra dimensions also reduces the probability [^10] Low intercommutation probability is typically assumed to be a property of superstrings. However, it has been argued that flux tubes of a non-abelian dark sector could have similar properties [71, 72]. for strings to scatter off each other, as they can now miss each other in the extra dimensions. If the strings were truly able to propagate within the extra-dimensional bulk, this would lower the reconnection probability by a factor of the volume, and quickly reach very small values. However, the position of the strings along the extra dimensions is a worldsheet scalar and corrections to e.g. the flat space metric will stabilize it. Using this logic, it has been argued [43] that warping effects (generically arising due to the presence of sources of energy along the extra dimensions) localize the strings in a small extra-dimensional region, yielding a reconnection probability of
$P\simeq 0.5g_{s}^{2}\left[\frac{6\pi}{\log\left(A\mathcal{V}\right)}\right]^{3}.$ (2.2)
@@
with $A$ a model-dependent, order one number (at weak warping). This depends very weakly in the volume. It is worth noting that brane-antibrane inflation, which naturally produces cosmic superstrings at its endpoint, has recently been reasessed [51, 52] and a potential with a hierarchical separation between the inflationary and late-time scales has been obtained. The volume modulus would evolve from its value during inflation to its exponentially large value at the true, late-time minimum, providing a concrete example of the ideas motivating this article.